6 more than one fourth of a number is two fifths of the number. Find the number.
Answer by Boreal(15235)

number =x
(x/4)+6=(2x/5)
multiply everything by 60
15x+360=24x
9x=360
x=40
(40/4)+6=16
40(2/5)=16
The number is 40.